Five years ago sister was four times older than Fido. Now she is only three times as old. How is old is Fido?
Answer by mananth(16946)

Fido - x years now
Sister --- 3x years
..
5 years ago
Fido'age = x-5 years
Sister's age = 3x-5 years
...
3x-5 =4(x-5)
3x-5=4x-20
4x-3x=20-5
x=15 Fido's age now
sister's age = 3x =45
...
m.ananth@hotmail.ca